# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

### v2

@@ 2.5 CART
Breiman et al. (1984)’s CART is a regression tree algorithm that grows a tree greedily in a top-down iterative fashion as follows: Starting with the entirely space {±$\{\pm 1\}^{d}$, each new cell $\mathcal{C}$ obtained in the algorithm is first checked for a stopping condition. If the condition is met, $\mathcal{C}$ is not split and is fixed as a leaf in the final tree. Otherwise, its split covariate is selected to be the maximizer of the impurity decrease, which is defined as
@@
When all leaves satisfy the stopping condition, the current tree $\mathfrak{T}$ is sometimes pruned to a subtree $\mathfrak{T}^{\prime}\subset\mathfrak{T}$. $\mathfrak{T}^{\prime}$ is then returned as the output to the algorithm. Several different stopping conditions and pruning rules are available (Breiman et al., 1984; Pedregosa et al., 2011). Our lower bounds will not depend on these choices so we will not go into further detail on what these are. Finally, for our upper bounds, we will make use the minimum impurity decrease stopping condition, which stops splitting $\mathcal{C}$ when it satisfies
$\frac{N(\mathcal{C})}{n}\max_{k\in[d]\backslash J(\mathcal{C})}\hat{\Delta}(k;\mathcal{C},\mathcal{D}_{n})<\gamma$ (14)
@@ 2.6 Random Forests
A random forest model is an ensemble of randomized CART trees. Specifically, for each CART tree, we fit it to a bootstrapped version of the dataset $\mathcal{D}_{n}$, denoted $\mathcal{D}_{n}^{*}$ and furthermore, at each iteration, a random subset $\mathcal{M}\subset[d]$ of some prespecified size is drawn uniformly at random, and the split covariate is selected from among these covariates, i.e., we set
$k(\mathcal{C})\coloneqq\operatorname*{arg\,max}_{k\in\mathcal{M}\backslash J(\mathcal{C})}\hat{\Delta}(k;\mathcal{C},\mathcal{D}_{n}^{*}).$ (15)
@@ 2.7 Greedy Tree Models
for a fixed function $\mathcal{F}$ that is invariant under the map $\left\{(X_{ik},Y_{i})\right\}_{i\in I(\mathcal{C})}\mapsto\left\{(-X_{ik},Y_{i})\right\}_{i\in I(\mathcal{C})}$. [^6] This condition is used in the proof of Lemma B.1. Since
$\mathcal{I}(\mathcal{T}_{k,z}(\mathcal{C});\mathcal{D}_{n})=\sum_{i\in I(\mathcal{C})}\left(\frac{|X_{i}^{k}+z|}{2}Y_{i}-\frac{\sum_{i\in I(\mathcal{C})}|X_{i}^{k}+z|Y_{i}}{\sum_{i\in I(\mathcal{C})}|X_{i}^{k}+z|}\right)^{2},$ (18)
for $z=-1,1$, we see that CART is indeed an instance of a greedy tree algorithm. We call an output of $\mathcal{A}$ a greedy tree model and denote it as $\hat{f}_{\operatorname{G}}$. Furthermore, if $\hat{f}_{\operatorname{G}}$ is the base learner in an ensemble constructed analogously to a random forest, we call the resulting ensemble a greedy tree ensemble and denote it as $\hat{f}_{\operatorname{G-E}}$.
@@ 3.2 Reduction to Covariate Selection Along Query Paths
$\begin{split}&\mathbb{E}_{\mathbf{X}}\left\{\left(\hat{f}(\mathbf{X};\mathcal{D}_{n},\Theta)-f^{*}(\mathbf{X})\right)^{2}~{}\vline~{}\mathbf{X}_{[d]\backslash S^{*}}=\mathbf{x}_{[d]\backslash S^{*}}\right\}\\ &=\operatorname{Var}\left\{f_{0}^{*}(\mathbf{X})\right\}+\left(\hat{f}(\mathbf{x};\mathcal{D}_{n},\Theta)-\mathbb{E}\left\{f_{0}^{*}(\mathbf{X})\right\}\right)^{2}\\ &\geq\operatorname{Var}\left\{f_{0}^{*}(\mathbf{X})\right\}.\end{split}$ (22)
@@
$\mathbb{E}_{\mathbf{X}\sim\nu}\left\{\mathbb{E}_{\mathbf{X}}\left\{\left(\hat{f}(\mathbf{X};\mathcal{D}_{n},\Theta)-f^{*}(\mathbf{X})\right)^{2}~{}\vline~{}\mathbf{X}_{[d]\backslash S^{*}}\right\}\mathbb{1}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}=\emptyset\right\}\right\}.$ (23)
@@
$\begin{split}&\mathbb{E}_{\mathbf{X},\Theta,\mathcal{D}_{n}}\left\{\left(\hat{f}(\mathbf{X};\mathcal{D}_{n},\Theta)-f^{*}(\mathbf{X})\right)^{2}\right\}\\ &\geq\mathbb{E}_{\mathbf{X}}\left\{\left(g(\mathbf{X})+h(\mathbf{X})-f^{*}(\mathbf{X})\right)^{2}\right\}\\ &=\mathbb{E}_{\mathbf{X}}\left\{\mathbb{E}_{\mathbf{X}}\left\{\left(g(\mathbf{X})+h(\mathbf{X})-f^{*}(\mathbf{X})\right)^{2}~{}\vline~{}\mathbf{X}_{[d]\backslash S^{*}}\right\}\right\}.\end{split}$ (26)
@@
$\begin{split}&\mathbb{E}_{\mathbf{X}}\left\{\left(g(\mathbf{X})+h(\mathbf{X})-f^{*}(\mathbf{X})\right)^{2}~{}\vline~{}\mathbf{X}_{[d]\backslash S^{*}}=\mathbf{x}_{[d]\backslash S^{*}}\right\}\\ &\geq\operatorname{Var}_{\mathbf{X}}\left\{h(\mathbf{X})-f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash S^{*}}=\mathbf{x}_{[d]\backslash S^{*}}\right\}\\ &\geq\operatorname{Var}\left\{f_{0}^{*}(\mathbf{X})\right\}-2\operatorname{Var}\left\{f_{0}^{*}(\mathbf{X})\right\}^{1/2}\operatorname{Var}_{\mathbf{X}}\left\{h(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash S^{*}}=\mathbf{x}_{[d]\backslash S^{*}}\right\}^{1/2}.\end{split}$ (27)
@@ 4 Greedy Trees and the XOR Function
Under Assumption 3.1, let $f_{0}^{*}$ be the XOR function. For any $10<\delta<$ 1, suppose $n\leq 2^{\delta(d-1)/2-2}$. Then the expected $L^{2}$ risk of any greedy tree model satisfies
$\mathfrak{R}\left(\hat{f}_{\operatorname{G}},f^{*}_{0},d,n\right)\geq(1-\delta)\operatorname{Var}\left\{f^{*}(\mathbf{x})\right\}.$ (29)
@@
$\mathfrak{R}\left(\hat{f}_{\operatorname{G-E}},f^{*}_{0},d,n\right)\geq(1-\kappa\delta)\operatorname{Var}\left\{f^{*}(\mathbf{x})\right\},$ (30)
@@
This implies that the response $Y$ is independent of the covariates in the modified dataset $\mathcal{D}_{n}^{(a)}$. Combining this with the fact that the covariates are drawn uniformly from {±$\{\pm 1\}^{d-1}$ as well as the symmetry of the splitting criterion, we see that the distribution of $\mathcal{D}_{n}^{(a)}$ is invariant to permutation of the covariate indices. As such, $J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)$ is the prefix of a uniformly random permutation of $[d]\backslash\{a\}$.
Step 3: Bounding covariate selection in coupled paths. For any $b\in[d]\backslash\{a\}$, we use the previous step to compute
@@
Taking a further expectation and using Lemma B.2 gives
@@
The technique of controlling a stochastic process by coupling it to another one which allows for more convenient calculations was also used by Tan and Vershynin (2023) to analyze the convergence of online SGD for the problem of phase retrieval.
@@ 5 Greedy Trees and Non-MSP Functions
It is clear that MSP is equivalent to $\mathcal{G}_{f^{*}}$ being a connected graph. Now let $\mathcal{S}_{\textsf{MSP}}$ denote the connected component of the empty set $\emptyset$ in $\mathcal{G}_{f^{*}}$ and set $S^{*}_{\textsf{MSP}}\coloneqq\cup_{S\in\mathcal{S}_{\textsf{MSP}}}S$. By adapting Lemma 4.2, we can show that any covariate $X_{k}$ with index not in $S^{*}_{\textsf{MSP}}$ is marginally undetectable, and thus rarely selected, on any query path of a greedy tree. We then modify Lemma 3.2 and Lemma 3.3 to translate lower bounds on the covariate selection probability into more nuanced expected risk lower bounds that can be stated in terms of the MSP residual of $f^{*}$, defined as $r_{\textsf{MSP}}\coloneqq\sum_{S\in\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}}\alpha_{S}\chi_{S}$ (see Lemma C.2 and Lemma C.3).
Naively following this recipe will produce a result with a sample dependence of $2^{\delta d/s+O(1)}$. In other words, we have a weaker lower bound on the necessary sample size as the support size of $f^{*}$ increases. This dependence does not always make sense: For instance, the bound obtained for the function $f^{*}\coloneqq\chi_{[d]}$ is trivial, even though this is in some sense the most complicated function possible. To overcome this limitation, we make a new definition. We say a subset $T\subset[d]\backslash S^{*}_{\textsf{MSP}}$ is a traversal of $\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}$ if for any $S\in\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}$, we have $|T\cap S|\geq 2$. It is easy to see that a traversal always exists. Furthermore, expected risk lower bounds can be established so long as we can derive lower bounds on the probability that the traversal is selected along a query path (see Lemma C.1). Putting the pieces together yields the first main result of our paper.
@@
Under Assumption 3.1, let $f_{0}^{*}=\sum_{S\in\mathcal{S}}\alpha_{S}\chi_{S}$ be any non-MSP function, and let $r_{\textsf{MSP}}\coloneqq\sum_{S\in\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}}\alpha_{S}\chi_{S}$ denote its MSP residual. Let $s_{\operatorname{T}}$ be the minimum size of a traversal for $\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}$ and let $s_{\textsf{MSP}}\coloneqq\left|S^{*}_{\textsf{MSP}}\right|$. For any $10<\delta<$ 1, suppose $\log_{2}n\leq\delta(d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1)/s_{\operatorname{T}}-2$. Then the expected $L^{2}$ risk of any greedy tree model satisfies
@@
It is possible to obtain similar lower bounds that hold with constant probability rather than in expectation. These results are available in Appendix I.
@@ 6 ERM Trees and Minimax Rates
Under Assumption 3.1, further suppose that $\varepsilon_{1},\varepsilon,\ldots,\varepsilon_{n}\sim_{\text{i.i.d.}}\mathcal{N}(0,\sigma^{2})$ and $2^{s}\log((d-s)/4)\geq 16\log 2$. Let $\mathfrak{F}_{s}$ denote all real-valued functions on {±$\left\{\pm 1\right\}^{s}$. Then we have the minimax lower bound
@@
Under Assumption 3.1, further suppose that $\varepsilon_{1},\varepsilon,\ldots,\varepsilon_{n}\sim_{\text{i.i.d.}}\mathcal{N}(0,\sigma^{2})$ and $\left\|f_{0}^{*}\right\|_{\infty}\leq M$. Consider the ERM tree model (11) with the constraint set $\Gamma$ comprising all tree structures of depth at most $s$. For any $f^{*}_{0}$, its expected $L^{2}$ risk satisfies the upper bound
@@ 7.1 Characterizing High-Dimensional Consistency for CART
When $f^{*}$ satisfies MSP, its associated hypergraph is connected. It is easy to verify that this implies that for any cell $\mathcal{C}$ on which $f^{*}$ is nonconstant, there exists a term $\alpha_{S}\chi_{S}$ in its Fourier expansion such that $S\backslash J(\mathcal{C})=\{j\}$ for some $j$. Unfortunately, this is not enough to guarantee that $X_{j}$ is marginally detectable, due to the possibility of cancellations. To see how these may arise, consider the function
@@
Fortunately, in some sense, exact cancellations happen rarely. To be more precise, first note that MSP is a condition on the Fourier support $\mathcal{S}$ of $f^{*}$ rather than on $f^{*}$ itself. Starting with such a collection, we see that the existence of a cancellation such as that in (44) is a linear constraint on the coefficients $\left\{\alpha_{S}\right\}_{S\in\mathcal{S}}$. If we draw these values from any distribution on $\mathbb{R}^{|\mathcal{S}|}$ that is absolutely continuous with respect to Lebesgue measure, the constraint hence forms a measure zero set, as does the union over all possible constraints. We say that a function $f^{*}$ satisfies the stable merged-staircase property (SMSP) if, for any $\mathcal{C}\subset\{\pm 1\}^{d}$, the restriction $f^{*}|_{\mathcal{C}}$ is MSP. We have thus shown that under a random perturbation of its coefficients, any MSP function is SMSP almost surely (see Proposition F.1).
Next, one can prove that $f^{*}$ is SMSP if and only if it satisfies the SID condition (7) for some value of $\lambda>0$ (see Proposition F.2). Previous work, specifically Chi et al. (2022), who first introduced the SID condition, as well as Mazumder and Wang (2024), provide upper bounds for the estimation error of CART given the SID condition. These results and proofs are stated for continuous covariates, but can easily be adapted to a discrete covariate setting. We revisit this last statement in the next subsection. Assuming this for now, our discussion thus far yields the following characterization theorem.
@@ 7.2 Sharp Upper Bounds for CART
The SID assumption does not always lead to sharp bounds. Consider for instance the $s$th-order AND function $f^{*}(\mathbf{x})=\mathbb{1}\left\{x_{j}=1\colon j=1,2,\ldots,s\right\}$. We may compute $\operatorname{Corr}^{2}\left\{f^{*}(\mathbf{X}),X_{j}\right\}=(2^{s}-1)^{-1}$, which implies that $f^{*}$ does not satisfy SID with any $\lambda$ below this value. Meanwhile, Theorem 2.3 in Mazumder and Wang (2024) gives the high probability deviation bound
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)=\tilde{O}\left(n^{-\phi(\lambda)}\right),$ (46)
where $\phi(\lambda)\coloneqq\frac{-\log_{2}(1-\lambda)}{1-\log_{2}(1-\lambda)}$. Plugging in $\lambda=(2^{s}-1)^{-1}$ shows that the bound is trivial unless $\log_{2}n=\tilde{\Omega}({2^{s}})$. We will soon show that this is unnecessarily pessimistic, with the true sample complexity of the order $\log_{2}n=2s+o(1)$.
To understand the source of pessimism in (46), note that the SID condition is used to guarantee that, for every cell $\mathcal{C}$, there exists a potential split that would decrease the residual variance within $\mathcal{C}$ by a constant factor of at most $1-\lambda$. Combined with concentration of impurity decrease values, one can then form a recursive inequality as follows: Letting $J(l)$ denote the squared bias of the CART tree grown to depth $l$, we have $J(l+1)\leq(1-\lambda(1+\epsilon)^{-1})J(l)$, where $\epsilon$ is an error that comes from finite sample randomness. The depth $l$ is then selected to optimize a bias-variance tradeoff. On the other hand, while the AND function has poor marginal signal strength at the root, the marginal signal strength improves as the tree gets deeper and more of the relevant covariates are selected. Indeed, we have $\operatorname{Corr}^{2}\left\{f^{*}(\mathbf{X}),X_{j}~{}|~{}X_{1}=X_{2}=\cdots=X_{j-1}=1\right\}=(2^{s-j}-1)^{-1}$. By only quantifying the worst case signal strength, the SID condition is unable to capture and exploit this nuance.
@@
$\max_{k\in[d]\backslash J(\mathcal{C})}\operatorname{Cov}^{2}\left\{f^{*}(\mathbf{X}),X_{k}~{}|~{}\mathbf{X}\in\mathcal{C}\right\}2^{-|J(\mathcal{C})\cap S^{*}|}\geq\lambda.$ (47)

### v3

@@ 2.5 CART
Breiman et al. (1984)’s CART is a regression tree algorithm that grows a tree greedily in a top-down iterative fashion as follows: Starting with the entire space $\{\pm 1\}^{d}$, each new cell $\mathcal{C}$ obtained in the algorithm is first checked for a stopping condition. If the condition is met, $\mathcal{C}$ is not split and is fixed as a leaf in the final tree. Otherwise, its split covariate is selected to be the maximizer of the impurity decrease, which is defined as
@@
When all leaves satisfy the stopping condition, the current tree $\mathfrak{T}$ is sometimes pruned to a subtree $\mathfrak{T}^{\prime}\subset\mathfrak{T}$. $\mathfrak{T}^{\prime}$ is then returned as the output to the algorithm. Several different stopping conditions and pruning rules are available (Breiman et al., 1984; Pedregosa et al., 2011). Our lower bounds will not depend on these choices so we will not go into further detail on what these are. Finally, for our upper bounds, we will make use of the minimum impurity decrease stopping condition, which stops splitting $\mathcal{C}$ when it satisfies
[…]
@@ 2.6 Random Forests
A random forest model is an ensemble of randomized CART trees. Specifically, for each CART tree, we fit it to a bootstrapped version of the dataset $\mathcal{D}_{n}$, denoted $\mathcal{D}_{n}^{*}$ and furthermore, at each iteration, a random subset $\mathcal{M}\subset[d]$ of some prespecified size is drawn uniformly at random with the split covariate is selected from among these covariates, i.e., we set
[…]
@@ 2.7 Greedy Tree Models
for a fixed function $\mathcal{F}$ that is invariant under the map $\left\{(X_{ik},Y_{i})\right\}_{i\in I(\mathcal{C})}\mapsto\left\{(-X_{ik},Y_{i})\right\}_{i\in I(\mathcal{C})}$. [^5] This condition is used in the proof of Lemma C.1. Since
$\mathcal{I}(\mathcal{T}_{k,z}(\mathcal{C});\mathcal{D}_{n})=\sum_{i\in I(\mathcal{C})}\left(\frac{|X_{ik}+z|}{2}Y_{i}-\frac{\sum_{i\in I(\mathcal{C})}|X_{ik}+z|Y_{i}}{\sum_{i\in I(\mathcal{C})}|X_{ik}+z|}\right)^{2},$ (18)
for z=1,1z=-$1,$ we see that CART is indeed an instance of a greedy tree algorithm. We call an output of $\mathcal{A}$ a greedy tree model and denote it as $\hat{f}_{\operatorname{G}}$. Furthermore, if $\hat{f}_{\operatorname{G}}$ is the base learner in an ensemble constructed analogously to a random forest, we call the resulting ensemble a greedy tree ensemble and denote it as $\hat{f}_{\operatorname{G-E}}$.
@@ 3.2 Reduction to Covariate Selection Along Query Paths
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4 Greedy Trees and the XOR Function
Under Assumption 3.1, let $f_{0}^{*}$ be the XOR function. For any $0<\delta<1$, suppose $n\leq 2^{\delta(d-1)/2-2}$. Then the expected $L^{2}$ risk of any greedy tree model satisfies
$\mathfrak{R}\left(\hat{f}_{\operatorname{G}},f^{*}_{0},d,n\right)\geq(1-\delta)\operatorname{Var}\left\{f_{0}^{*}(\mathbf{X})\right\}.$ (29)
@@
$\mathfrak{R}\left(\hat{f}_{\operatorname{G-E}},f^{*}_{0},d,n\right)\geq(1-\kappa\delta)\operatorname{Var}\left\{f_{0}^{*}(\mathbf{X})\right\},$ (30)
@@
[…]
@@
Taking a further expectation and using Lemma C.2 gives
@@
The technique of controlling a stochastic process by coupling it to another one which allows for more convenient calculations was also used by Tan and Vershynin (2023) to analyze the convergence of online SGD for the problem of phase retrieval, and by Abbe and Sandon (2023); Abbe and Boix-Adserà (2022) to analyze gradient descent on leap functions.
@@ 5 Greedy Trees and Non-MSP Functions
It is clear that MSP is equivalent to $\mathcal{G}_{f^{*}}$ being a connected graph. Now let $\mathcal{S}_{\textsf{MSP}}$ denote the connected component of the empty set $\emptyset$ in $\mathcal{G}_{f^{*}}$ and set $S^{*}_{\textsf{MSP}}\coloneqq\cup_{S\in\mathcal{S}_{\textsf{MSP}}}S$. By adapting Lemma 4.2, we can show that any covariate $X_{k}$ with index not in $S^{*}_{\textsf{MSP}}$ is marginally undetectable, and thus rarely selected, on any query path of a greedy tree. We then modify Lemma 3.2 and Lemma 3.3 to translate lower bounds on the covariate selection probability into more nuanced expected risk lower bounds that can be stated in terms of the MSP residual of $f^{*}$, defined as $r_{\textsf{MSP}}\coloneqq\sum_{S\in\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}}\alpha_{S}\chi_{S}$ (see Lemma D.2 and Lemma D.3).
[…]
@@
Under Assumption 3.1, let $f_{0}^{*}=\sum_{S\in\mathcal{S}}\alpha_{S}\chi_{S}$ be any non-MSP function, and let $r_{\textsf{MSP}}\coloneqq\sum_{S\in\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}}\alpha_{S}\chi_{S}$ denote its MSP residual. Let $s_{\operatorname{T}}$ be the minimum size of a traversal for $\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}$ and let $s_{\textsf{MSP}}\coloneqq\left|S^{*}_{\textsf{MSP}}\right|$. For any $0<\delta<1$, suppose $\log_{2}n\leq\delta(d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1)/s_{\operatorname{T}}-2$. Then the expected $L^{2}$ risk of any greedy tree model satisfies
@@
It is possible to obtain similar lower bounds that hold with constant probability rather than in expectation. These results are available in Appendix J.
@@ 6 ERM Trees and Minimax Rates
Under Assumption 3.1, further suppose that $\varepsilon_{1},\varepsilon,\ldots,\varepsilon_{n}\sim_{\text{i.i.d.}}\mathcal{N}(0,\sigma^{2})$ and $2^{s}\log((d-s)/4)\geq 16\log 2$. Let $\mathfrak{F}_{s}$ denote all real-valued functions on $\left\{\pm 1\right\}^{s}$. Then we have the minimax lower bound
@@
Under Assumption 3.1, further suppose that $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}\sim_{\text{i.i.d.}}\mathcal{N}(0,\sigma^{2})$ and $\left\|f_{0}^{*}\right\|_{\infty}\leq M$. Consider the ERM tree model (11) with the constraint set $\Gamma$ comprising all tree structures of depth at most $s$. For any $f^{*}_{0}$, its expected $L^{2}$ risk satisfies the upper bound
@@ 7.1 Characterizing High-Dimensional Consistency for CART
[…]
@@
Fortunately, in some sense, exact cancellations happen rarely. To be more precise, first note that MSP is a condition on the Fourier support $\mathcal{S}$ of $f^{*}$ rather than on $f^{*}$ itself. Starting with such a collection, we see that the existence of a cancellation such as that in (44) is a linear constraint on the coefficients $\left\{\alpha_{S}\right\}_{S\in\mathcal{S}}$. If we draw these values from any distribution on $\mathbb{R}^{|\mathcal{S}|}$ that is absolutely continuous with respect to Lebesgue measure, the constraint hence forms a measure zero set, as does the union over all possible constraints. Thus, our results may equivalently be read deterministically for any fixed set of coefficients outside this exceptional set.
We say that a function $f^{*}$ satisfies the stable merged-staircase property (SMSP) if, for any $\mathcal{C}\subset\{\pm 1\}^{d}$, the restriction $f^{*}|_{\mathcal{C}}$ is MSP. We have thus shown that under a random perturbation of its coefficients, any MSP function is SMSP almost surely (see Proposition G.1).
Next, one can prove that $f^{*}$ is SMSP if and only if it satisfies the SID condition (7) for some value of $\lambda>0$ (see Proposition G.2). Previous work, specifically Chi et al. (2022), who first introduced the SID condition, as well as Mazumder and Wang (2024), provide upper bounds for the estimation error of CART given the SID condition. These results and proofs are stated for continuous covariates, but can easily be adapted to a discrete covariate setting. We revisit this last statement in the next subsection. Assuming this for now, our discussion thus far yields the following characterization theorem.
@@ 7.2 Sharp Upper Bounds for CART
The SID assumption does not always lead to sharp bounds. Consider for instance the $s$th-order AND function $f^{*}(\mathbf{x})=\mathbb{1}\left\{x_{j}=1\colon j=1,2,\ldots,s\right\}$. We may compute $\operatorname{Corr}^{2}\left\{f^{*}(\mathbf{X}),X_{j}\right\}=(2^{s}-1)^{-1}$, which implies that $f^{*}$ does not satisfy SID with any $\lambda$ below this value. Note that the SID condition is used to guarantee that, for every cell $\mathcal{C}$, there exists a potential split that would decrease the residual variance within $\mathcal{C}$ by a constant factor of at most $1-\lambda$. Combined with concentration of impurity decrease values, one can then form a recursive inequality of the form $J(l+1)\leq(1-\lambda/(1+\epsilon))J(l)$, where $J(l)$ denotes the squared bias of the CART tree grown to depth $l$ and $\epsilon$ is a finite-sample error. Iterating gives $J(l)\leq J(0)\exp(-\lambda l/(1+\epsilon))$, while the variance at depth $l$ is $O(2^{l}/n)$, up to poly-log factors in $n$ and $d$. Balancing bias and variance requires $\exp(-\lambda l/(1+\epsilon))\approx 2^{l}/n$, i.e., $l\approx\log n/(\lambda/(1+\epsilon)+\log 2)$, in which case we obtain the optimized risk bound $O(n^{-\lambda/((1+\epsilon)\log 2+\lambda)})$. Substituting $\lambda=(2^{s}-1)^{-1}$ shows that the exponent $\lambda/((1+\epsilon)\log 2+\lambda)$ is of the order $2^{-s}$, so the bound is trivial unless $\log_{2}n=\Omega(2^{s})$. We will soon show that this is unnecessarily pessimistic, with the true sample complexity of the order $\log_{2}n=2s+o(1)$.
On the other hand, while the AND function has poor marginal signal strength at the root, the marginal signal strength improves as the tree gets deeper and more of the relevant covariates are selected. Indeed, we have $\operatorname{Corr}^{2}\left\{f^{*}(\mathbf{X}),X_{j}~|~X_{1}=X_{2}=\cdots=X_{j-1}=1\right\}=(2^{s-j}-1)^{-1}$. By only quantifying the worst case signal strength, the SID condition is unable to capture and exploit this nuance.
@@
[…]